OpenStudy (anonymous):

what is the domain of g(k) = 2(k^2)+4k-6

OpenStudy (anonymous):

All real numbers (it has no denominator, so it can never be undefined)

OpenStudy (anonymous):

how do you right that with correct set notations?

OpenStudy (anonymous):

\[(-\infty, \infty)\]

OpenStudy (anonymous):

do you mind helping me with my other problems?

OpenStudy (anonymous):

Just post them

OpenStudy (anonymous):

domain of a(r) = r + 1/(r-1)

OpenStudy (anonymous):

domain of a(r) = r + 1/(r-1)